# Rates of Convergence in the Central Limit Theorem for Markov Chains, with an Application to TD Learning

### v2

@@ 1 Introduction
TD learning is the most widely studied algorithm for evaluating the performance of a given policy from data in Markov Decision Processes [1]. The convergence of TD learning with decaying stepsizes was proved in [2]. Over the last few years, there has been a resurgent interest in understanding the non-asymptotic convergence behavior of TD learning and further using it to study other other reinforcement algorithms which use TD learning within their framework. The first result of this type was obtained in [3] who used a projection step in their version of the TD learning algorithm. The first finite-time bounds for unprojected TD learning were obtained in [4], who studied both fixed step-sizes and decaying step-sizes. TD learning can be viewed as a special case of linear stochastic approximation with Markovian and multiplicative noise and as a special case of Markovian jump linear systems [5]. The results in [4] were extended to nonlinear, contractive stochastic approximation algorithms in [6], who also derived finite-time performance bounds for specific choices of decaying step-size rules, which also allowed them to study the robustness of different step-size choices. Finite-time performance bounds for variants of TD learning and other schemes have been studied in [7, 8, 9] and concentration results for TD learning have been obtained in [10, 11].
While most of the above papers study standard TD learning, in practice, it is well-known that averaging the iterates can help improve the asymptotic variance. This idea was originally studied for stochastic approximation and stochastic gradient descent in the seminal works in [12, 13, 14]; see [15] for a finite-time analysis. More recently, the asymptotic behavior of TD learning and other algorithms with averaging has been studied in [16, 17, 18]; see also [19] for related work. As in the papers on stochastic approximation, the conclusion in [16] is that, for a suitable choice of decaying stepsizes, averaging the output leads to efficient estimators in the sense that the variance is minimized. Further, [16] also established functional central limit theorems and central limit theorems for TD learning and other reinforcement learning algorithms. The work in [17] shows order-optimal rate of convergence for the mean-squared error, while [18] obtains higher moment bounds and high probability deviation bounds. In contemporaneous work to this paper, finite-time performance bounds on the covariance of the averaged TD learning iterates have been obtained in [20] using more robust step-size rules. Their model is more general; they study two-time scale linear stochastic approximation algorithms, with TD learning with averaging being a special case of the models that they study. A central limit theorem for two time-scale stochastic approximation has also been established for two time-scale stochastic approximation in [21]. Related work also includes a central limit theorem for Polyak-Ruppert averaged Q-learning [22].
@@
Our goal in this paper is to obtain bounds on the rate of convergence to normality to complement the central limit theorem in [16]. To this end, we first establish a result for martingale differences by combining the Lindeberg decomposition along with Stein’s method, which was first used in [23] for scalar valued martingale differences. This key idea in [23] was extended to the case of vector-valued martingales in [24] who also used it to obtain a non-asymptotic version of the central limit theorem in [14]. Our approach builds upon the ideas in [23, 24], but we also introduce new techniques in order to obtain stronger results and to deal with the differences in our model compare to [24]:
@@
  • We would like our results to hold when the deviation from normality is measured using the Wasserstein distance. The results in [24] provide bounds only for weaker distances. Therefore, we have to obtain results for more general test functions than the ones considered in [24]. To do so, we leverage the results on the regularity of solutions to Stein’s equation for vector-valued random variables in [25, 26], which makes our analysis considerably different.
@@
  • To obtain rates of convergence for the Markov chain central limit theorem, we convert a Markov chain into a martingale difference noise using Poisson’s equation; see [27, 28, 29], for example. To the best of our knowledge, this idea has not been used to derive rates of convergence for the central limit theorem for functions of Markov chains.
  • In addition to the Markovian nature of our noise, we have to deal with the fact that the noise multiplies the TD learning parameters, unlike the case of stochastic gradient descent, where the martingale difference noise is additive.
@@
A notable difference between our results and the results in [23, 24] for martingales is the fact that we explicitly relate our convergence bounds to the rate of convergence of conditional covariances of martingale differences to the asymptotic covariance (see item (2) in Assumption 2 and Theorem 1). This is important to us since our goal is relate our results on TD learning to the asymptotic covariance results in [16]. Therefore, we cannot use the assumptions on the variance in [23, Theorem 1] or the covariance assumption in [24, Equation 3.1]. It is also worth remarking that Stein’s method to derive rates of convergence for the martingale central limit theorem is relatively new. The early papers on obtaining rates of convergence for the martingale central limit theorem use other techniques; see, for example, [30] and other references in [23]. Here, we use Stein’s method because we found it easier to quantify the impact of the asymptotic variance on the rate of convergence using Stein’s method. Specifically, we show that the regularity results in [25, 26] and the calculations in [23, 25] can be used, along with insights from Markov chain theory, to understand the impact of the asymptotic covariance term on the rates of convergence in the martingale and Markov chain central limit theorems.
@@
  • We derive non-asymptotic central limit theorems for martingales and Markov chains by exploiting ideas from [23, 25, 26]. One key difference compared to [23, 24] is that we write our results explicitly in terms of the asymptotic covariance of the martingale, which is useful when we study Markov chains.
  • We then extend the results to Markov chains using Poisson’s equation to relate Markov chains to martingales. In this case, the asymptotic covariance can be characterized as in well known in the literature; see [29], for example. The error bound on the deviation from normality in the case of martingale difference sequences appears to be complicated. Interestingly however, under appropriate assumptions, the expressions simplify remarkably in the case of Markov chains and leads to a simple convergence rate result.
  • Finally, we apply the Markov chain results, along with the ideas in [14], to estimate the rate of convergence in the central limit theorem for TD learning with Polyak-Ruppert averaging. An advantage of the approach here is that our result directly establishes a central limit theorem for the scaled averaged of the iterates, instead of first establishing a function central limit theorem for the entire trajectory of the iterates and then establishing the central limit theorem.
It is important to note that, unlike [14], our result only holds for averaging along with an appropriate decaying step-size rule. Specifically, the result does not hold for fixed step-size TD learning with averaging due to the bias issue pointed out in [16, 31, 32]); also [33, 34, 35] for challenges in dealing with Markov noise in other problems. A nonasymptotic central limit theorem has been established for constant step-size stochastic gradient descent in [36], but the model considered here is different from the problem in [36] (for example, Markovian noise, constant step-size vs decaying step-size) and therefore, the techniques there do not seem to apply to our problem.
@@ 2.1 Preliminaries
    $\Sigma_{k}=E(m_{k}m_{k}^{T}|\mathcal{F}_{k-1})$ exists and further assume that
    $\lim_{n\rightarrow\infty}\frac{\Sigma_{1}+\ldots+\Sigma_{n}}{n}=\Sigma_{\infty}$
    almost surely for some $\Sigma_{\infty}\succ 0,$ where $\succ 0$ indicates positive definiteness.
@@
Item (1) in the assumption is motivated by the fact that we want to extend the techniques in [25] to martingales and item (2) is motivated by the fact that one can show that $\Sigma_{\infty}$ exists for certain martingales derived from some Markov chains, as we will see in the next section.
@@
converges to $N(0,\Sigma_{\infty})$, and if so, we are interested in the rate of convergence. We will use Stein’s method to establish such a convergence and to estimate the rate of convergence [37, 38, 39].
@@
The following results from [25, Propositions 2.2 and 2.3] and [26, Theorem 3.1] will be useful to us and so we present them together as a lemma for future reference.
@@
where $\Delta$ and normal-$\nabla$ are the Laplacian and gradient operators, respectively, and $Z\sim N(0,I)$. Then, $f$ has the following regularity properties for any $\beta\in(0,1)$:
@@
Next, we state a well-known result regarding Ornstein-Uhlenbeck processes which will be useful later to establish the nonasymptotic version of the martingale central limit theorem (for example, the result can be easily inferred from the discussions on generators and linear stochastic differential equations in [40]).
@@
Thus, if $\tilde{Z}_{k}\sim N(0,\Sigma),$ then $E(\mathcal{A}f(\tilde{Z}_{k}))=0$ for any twice differentiable function $f.$ normal-$\diamond$
It is worth noting that Stein’s equation (2) can be motivated by the above lemma when $A=-I$ and $BB^{T}=2I,$ in which case the right-hand side of (2) is the same as $\mathcal{A}f(u)$. Thus, if $u$ in (2) is replaced by a random vector, then the left-hand side of (2) is a measure of the difference between the distribution of that random vector and a standard Gaussian distribution while the right-hand side will hopefully be small if the distribution of the random vector is close to a standard Gaussian distribution (see, for example, [41, 42, 43]).
@@ 2.2 Rate of Convergence
where $T_{k}=\Sigma_{\infty}^{1/2}(Z_{k}+\ldots+Z_{n}),$ where $Z_{i}$ are i.i.d. $N(0,I)$ random vectors, independent of the martingale difference sequence. Then, using the Lindeberg decomposition [23], we have
@@
The rest of the proof uses the vector extensions of the techniques used in [23] for scalar random variables. As in well known in the Stein’s method literature, going from scalar-valued random variables to random vectors is non-trivial. To deal with random vectors, we use the ideas in [25, Theorem 3.1] which were developed for i.i.d. random variables. One of our contributions is to show that these ideas can be used for martingales using the Lindeberg decomposition. While doing so, we also have to deal with the asymptotic covariance $\Sigma_{\infty},$ which does not appear in [23, 24] due to their assumptions, but will be important to us when we study Markov chains.
@@
where $\tilde{Z}_{k}\sim N(0,\sqrt{\frac{n-k}{n-k+1}}I)$ has the same distribution as $\frac{1}{\sqrt{n-k+1}}\Sigma_{\infty}^{-1/2}T_{k+1}$. Next, we rewrite the expression as a sum of four terms as follows:
@@
The above theorem is useful if one can show that the upper bound on the Wasserstein distance goes to zero as $n\rightarrow\infty,$ in which case, the rate at which the upper bound goes to zero provides some information on the rate of convergence. In the next section, we will show that such a result can be obtained for a certain martingale derived from the solution of Poisson’s equation for Markov chains and obtain a bound on the rate of convergence for the central limit theorem for functions of Markov chains.
@@ 3 Markov Chain Central Limit Theorem
Consider a time-homogeneous Markov chain $\{X_{k}\}_{k\geq 0}$ on a state-space $\mathcal{S}$ with some initial distribution $\eta.$ For simplicity, we will assume that the state space is discrete, although that assumption may be relaxed by using the theory for general state spaces in [29, 44] for some of our results. Let $r(X_{k})\in\Re^{d}$ denote a vector of rewards accrued when in state $X_{k}.$ We are interested in rates of convergence to normality of
@@
As is common in the literature, we will study the above scaled sum by relating the reward process to a martingale via Poisson’s equation [29, 45, 46, 47].
@@
We make the following assumptions about the Markov chain.
@@
    The Markov chain has a unique stationary distribution $\pi$ and is uniformly geometrically ergodic, i.e., there exists $K_{1}>0$ and $\rho\in[0,1)$ such that $\forall x\in\mathcal{S},$ we have
@@
    where $\bar{r}=E_{y\sim\pi}(r(y)),$ such that $||V(X_{k})||\leq M_{V},\forall k\geq 0$ for some $M_{V}<\infty.$
@@
    Define $m_{k}=V(X_{k})-E(V(X_{k})|X_{k-1})$ for $k\geq 1.$ Clearly, $E(m_{k}|\mathcal{F}_{k-1})=0.$ Further, we assume that $\{m_{k}\}_{k\geq 1}$ is a martingale difference sequence with respect to the filtration $\{\mathcal{F}_{k}\}_{k\geq 0}$ satisfying Assumption 1 item (2), with $\mathcal{F}_{k}=\sigma(X_{0},X_{1},X_{2},\ldots,X_{k})$ and $\Sigma_{\infty}$ given by
@@
    where $P_{ij}$ is the $(i,j)^{\rm th}$ element of the probability transition matrix.
$\diamond$
Clearly, under Assumption 2 item (2), there exists constants $M_{2+\beta}$ and $M_{\beta}$ such that
Assumption 2 item (2) on the solution to Poisson’s equation is made for simplicity. It can be weakened considerably as one can see by following the proof of the subsequent theorem. The assumption here would be sufficient for our application to TD learning in the next section, so we choose to make this assumption.
@@
$d_{\mathcal{W}}(U_{n},\Sigma_{\infty}^{1/2}Z)=\mathcal{O}\Bigl{(}\frac{1}{n^{\beta/2}}\Bigr{)},$
@@
We start with the following standard use of Poisson’s equation and the definition of $m_{k}$ in Assumption 2 (see, for example, [29]) to write
@@
$U_{n}=W_{n}+\frac{V(X_{0})-V(X_{n})}{\sqrt{n}},\quad\mathrm{where}\quad W_{n}=\frac{1}{\sqrt{n}}\sum_{k=1}^{n}m_{k}.$
@@
By our assumption of uniform geometric ergodicity and Assumption 2 item (2), the above expression goes to zero geometrically fast and thus,
@@
where $K_{3}$ is a problem-dependent constant. ∎
Our proof relied on assumptions about the solution to Poisson’s equation (5). The Poisson equation is the same as the average-reward relative value function equation, and there are simple conditions for the existence of a solution for this equation in the case of finite state-space Markov chains [48]. Examples of models with more general state-spaces where one can obtain solutions to Poisson’s equation satisying our assumptions can be found in [49].
As in [25], we can obtain the following corollary by choosing $\beta=1-2/\log n$ and noting that $\tilde{C}_{1}$ and $\tilde{C}_{2}$ become $O(\log n)$ in that case.
Corollary 1.
Under Assumption 3, the following rate of convergence holds:
$d_{\mathcal{W}}(\sqrt{n}(U_{n},\Sigma_{\infty}^{1/2}Z)=\mathcal{O}\Bigl{(}\frac{\log n}{\sqrt{n}}\Bigr{)},$
where $\Sigma_{\infty}$ has been defined in Assumption 2, with $r(X_{k})$ as in Assumption 3. normal-$\diamond$

### v3

@@ 1 Introduction
Starting with the seminal work of (Polyak and Juditsky 1992), where they analyzed the averaging rule of (Polyak 1990, Ruppert 1988), the central limit theorem for martingales has been a valuable tool to understand asymptotic efficiency of stochastic approximation algorithms for optimization. More recently, the central limit theorem for functions of Markov chains has also been used to study the asymptotic efficiency of machine learning algorithms (Borkar et al. 2021, Hu et al. 2024). However, in machine learning applications, finite-time (rather than asymptotic) bounds are preferred to understand the sample complexity of algorithms. However, to the best of our knowledge, there is limited work on the central limit theorems for vector-valued martingales and vector-valued functions of Markov chains that would allow one to both quantify asymptotic efficiency and obtain finite-time bounds for learning and optimization algorithms. A recent paper by (Anastasiou et al. 2019) derives a rate of convergence for the central limit theorem vector-valued martingales; however, the notion of distance between probability distributions in that paper is not sufficient for some applications.
@@
Our goal in this paper is to obtain bounds on the rate of convergence to normality to complement the central limit theorem for Markov chains (Meyn and Tweedie 2012). To this end, we first establish a result for martingale differences by combining the Lindeberg decomposition along with Stein’s method, which was first used in (Röllin 2018) for scalar valued martingale differences. This key idea in (Röllin 2018) was extended to the case of vector-valued martingales in (Anastasiou et al. 2019) who also used it to obtain a non-asymptotic version of the central limit theorem in (Polyak and Juditsky 1992). Our approach builds upon the ideas in (Röllin 2018, Anastasiou et al. 2019), but we also introduce new techniques in order to obtain stronger results:
@@
  • We would like our results to hold when the deviation from normality is measured using the Wasserstein distance. The results in (Anastasiou et al. 2019) provide bounds only for weaker distances. Therefore, we have to obtain results for more general test functions than the ones considered in (Anastasiou et al. 2019). To do so, we leverage the results on the regularity of solutions to Stein’s equation for vector-valued random variables in (Gallouët et al. 2018, Fang et al. 2019), which makes our analysis considerably different.
@@
  • To obtain rates of convergence for the Markov chain central limit theorem, we convert a Markov chain into a martingale difference noise using Poisson’s equation; see (Metivier and Priouret 1984, Benveniste et al. 2012, Douc et al. 2018), for example. To the best of our knowledge, this idea has not been used to derive convergence rates for the central limit theorem for functions of Markov chains.
@@
A difference between our results and the results in (Röllin 2018, Anastasiou et al. 2019) for martingales is the fact that we state our convergence bounds in terms of a positive definite matrix $\Sigma_{\infty}$ which can be viewed as the limit of conditional covariances of martingale differences (see Theorem 1 and the short discussion after the theorem). This is important to us since our goal is to state our results for Markov chains in terms of the asymptotic covariance, which is important in applications; see (Borkar et al. 2021), for example. Therefore, our assumptions are stated differently than in (Röllin 2018, Theorem 1) or in (Anastasiou et al. 2019, Equation 3.1). It is also worth remarking that Stein’s method to derive rates of convergence for the martingale central limit theorem is relatively new. The early papers on obtaining rates of convergence for the martingale central limit theorem use other techniques; see, for example, (Bolthausen 1982) and other references in (Röllin 2018). Here, we use Stein’s method because we found it easier to quantify the impact of the asymptotic variance on the rate of convergence using Stein’s method. Specifically, we show that the regularity results in (Gallouët et al. 2018, Fang et al. 2019) and the calculations in (Röllin 2018, Gallouët et al. 2018) can be used, along with insights from Markov chain theory, to understand the impact of the asymptotic covariance term on the rates of convergence in the martingale and Markov chain central limit theorems.
As an application, we apply the Markov chain central limit theorem to study the distribution of error in TD learning. TD learning is the most widely studied algorithm for evaluating the performance of a given policy from data in Markov Decision Processes (Sutton and Barto 2018). The convergence of TD learning with decaying stepsizes was proved in (Tsitsiklis and Van Roy 1996). Over the last few years, there has been a resurgent interest in understanding the non-asymptotic convergence behavior of TD learning and further using it to study other reinforcement algorithms which use TD learning within their framework. The first result of this type was obtained in (Bhandari et al. 2018) who used a projection step in their version of the TD learning algorithm. The first finite-time bounds for unprojected TD learning were obtained in (Srikant and Ying 2019), who studied both fixed step-sizes and decaying step-sizes. TD learning can be viewed as a special case of linear stochastic approximation with Markovian and multiplicative noise and as a special case of Markovian jump linear systems (Hu and Syed 2019). The results in (Srikant and Ying 2019) were extended to nonlinear, contractive stochastic approximation algorithms in (Chen et al. 2023a), who also derived finite-time performance bounds for specific choices of decaying step-size rules, which also allowed them to study the robustness of different step-size choices. Finite-time performance bounds for variants of TD learning and other schemes have been studied in (Gupta et al. 2019, Doan 2021, 2022) and concentration results for TD learning have been obtained in (Telgarsky 2022, Chen et al. 2023b).
While most of the above papers study standard TD learning, in practice, it is well-known that averaging the iterates can help improve the asymptotic variance. This idea was originally studied for stochastic approximation and stochastic gradient descent in the seminal works in (Ruppert 1988, Polyak 1990, Polyak and Juditsky 1992); see (Moulines and Bach 2011) for a finite-time analysis. More recently, the asymptotic behavior of TD learning and other algorithms with averaging has been studied in (Borkar et al. 2021, Mou et al. 2021, Durmus et al. 2022); see also (Li et al. 2023a) for related work. As in the papers on stochastic approximation, the conclusion in (Borkar et al. 2021) is that, for a suitable choice of decaying stepsizes, averaging the output leads to efficient estimators in the sense that the variance is minimized. Further, (Borkar et al. 2021) also established functional central limit theorems and central limit theorems for TD learning and other reinforcement learning algorithms. The work in (Mou et al. 2021) shows order-optimal rate of convergence for the mean-squared error, while (Durmus et al. 2022) obtains higher moment bounds and high probability deviation bounds. In contemporaneous work to this paper, finite-time performance bounds on the covariance of the averaged TD learning iterates have been obtained in (Haque et al. 2023) using more robust step-size rules. Their model is more general; they study two-time scale linear stochastic approximation algorithms, with TD learning with averaging being a special case of the models that they study. A central limit theorem for two time-scale stochastic approximation has also been established for two time-scale stochastic approximation in (Hu et al. 2024). Related work also includes a central limit theorem for Polyak-Ruppert averaged Q-learning (Li et al. 2023b).
@@
  • We derive non-asymptotic central limit theorems for martingales and Markov chains by exploiting ideas from (Röllin 2018, Gallouët et al. 2018, Fang et al. 2019).
  • We then extend the results to Markov chains using Poisson’s equation to relate Markov chains to martingales. In this case, the asymptotic covariance can be characterized as is well known in the literature; see (Douc et al. 2018), for example.
  • Finally, we apply the Markov chain results, along with the ideas in (Polyak and Juditsky 1992), to estimate the rate of convergence in the central limit theorem for TD learning with Polyak-Ruppert averaging. An advantage of the approach here is that our result directly establishes a central limit theorem for the scaled averaged of the iterates, instead of first establishing a function central limit theorem for the entire trajectory of the iterates and then establishing the central limit theorem.
  • In addition to the difference in the notion of distance considered here and in (Anastasiou et al. 2019), in the TD learning application, we have to deal with the fact that the noise multiplies the TD learning parameters, whereas the noise is additive in the stochastic gradient application considered in (Anastasiou et al. 2019).
It is important to note that, unlike (Polyak and Juditsky 1992), our result only holds for averaging along with an appropriate decaying step-size rule. Specifically, the result does not hold for fixed step-size TD learning with averaging due to the bias issue pointed out in (Borkar et al. 2021, Lauand and Meyn 2023, Huo et al. 2023)); also see (Nagaraj et al. 2020, Roy and Balasubramanian 2023, Jain et al. 2018) for challenges in dealing with Markov noise in other problems. A nonasymptotic central limit theorem has been established for constant step-size stochastic gradient descent in (Dieuleveut et al. 2020), but the model considered here is different from the problem in (Dieuleveut et al. 2020) (for example, Markovian noise, constant step-size vs decaying step-size) and therefore, the techniques there do not seem to apply to our problem.
@@ 2.1 Preliminaries
    $\Sigma_{k}=E(m_{k}m_{k}^{T}|\mathcal{F}_{k-1})$ exists for all $k$.
@@
Item (1) in the assumption is motivated by the fact that we want to extend the techniques in (Gallouët et al. 2018) to martingales .
@@
converges to $N(0,\Sigma_{\infty})$, and if so, we are interested in the rate of convergence. We will use Stein’s method to establish such a convergence and to estimate the rate of convergence (Stein 1972, Ross 2011, Chatterjee 2014).
@@
The following results from (Gallouët et al. 2018, Propositions 2.2 and 2.3) and (Fang et al. 2019, Theorem 3.1) will be useful to us and so we present them together as a lemma for future reference.
@@
where $\Delta$ and $\nabla$ are the Laplacian and gradient operators, respectively, and $Z\sim N(0,I)$. Then, $f$ has the following regularity properties for any $\beta\in(0,1)$:
@@
Next, we state a well-known result regarding Ornstein-Uhlenbeck processes which will be useful later to establish the nonasymptotic version of the martingale central limit theorem (for example, the result can be easily inferred from the discussions on generators and linear stochastic differential equations in (Arnold 1974)).
@@
Thus, if $\tilde{Z}_{k}\sim N(0,\Sigma),$ then $E(\mathcal{A}f(\tilde{Z}_{k}))=0$ for any twice differentiable function $f.$diamond
It is worth noting that Stein’s equation (2) can be motivated by the above lemma when $A=-I$ and $BB^{T}=2I,$ in which case the right-hand side of (2) is the same as $\mathcal{A}f(u)$. Thus, if $u$ in (2) is replaced by a random vector, then the left-hand side of (2) is a measure of the difference between the distribution of that random vector and a standard Gaussian distribution while the right-hand side will hopefully be small if the distribution of the random vector is close to a standard Gaussian distribution (see, for example, (Barbour 1990, Gotze 1991, Chen et al. 2010)).
@@ 2.2 Rate of Convergence
where $T_{k}=\Sigma_{\infty}^{1/2}(Z_{k}+\ldots+Z_{n}),$ where $Z_{i}$ are i.i.d. $N(0,I)$ random vectors, independent of the martingale difference sequence. Then, using the Lindeberg decomposition (Röllin 2018), we have
@@
The rest of the proof uses the vector extensions of the techniques used in (Röllin 2018) for scalar random variables. As in well known in the Stein’s method literature, going from scalar-valued random variables to random vectors is non-trivial. To deal with random vectors, we use the ideas in (Gallouët et al. 2018, Theorem 3.1) which were developed for i.i.d. random variables. One of our contributions is to show that these ideas can be used for martingales using the Lindeberg decomposition. While doing so, we also have to deal with the asymptotic covariance $\Sigma_{\infty},$ which does not appear in (Röllin 2018, Anastasiou et al. 2019) due to their assumptions, but will be important to us when we study Markov chains.
@@
where $\tilde{Z}_{k}\sim N(0,\sqrt{\frac{n-k}{n-k+1}}I)$ has the same distribution as $\frac{1}{\sqrt{n-k+1}}\Sigma_{\infty}^{-1/2}T_{k+1}$ and is independent of all other random variables . Next, we rewrite the expression as a sum of four terms as follows:
@@
$\Box$
Note that if $\Sigma_{\infty}$ is such that $\Sigma_{k}\rightarrow\Sigma_{\infty}$ as $k\rightarrow\infty,$ then the rate at which this convergence occurs will determine the rate of convergence in the martingale CLT. One can characterize such a rate explicitly in the case of Markov chains as we will see in the next section.
@@ 3 Markov Chain Central Limit Theorem
Consider a time-homogeneous Markov chain $\{X_{k}\}_{k\geq 0}$ on a state space $\mathcal{S}$ with some initial distribution $\eta.$ Let $r(X_{k})\in\Re^{d}$ denote a vector of rewards accrued when in state $X_{k}.$ We are interested in rates of convergence to normality of
@@
As is common in the literature, we will study the above scaled sum by relating the reward process to a martingale via Poisson’s equation (Douc et al. 2018, Makowski and Shwartz 2002, Glynn and Meyn 1996, Glynn and Infanger 2023).
3.1 Finite State Space Markov Chains
In this subsection, we assume that $\mathcal{S}$ is finite and extend the results to more general state spaces in the next subsection.
@@ 3.1 Finite State Space Markov Chains
We assume that the Markov chain takes values in finite set $\mathcal{S}$ and is irreducible and aperiodic. $\diamond$
The assumption has a few well-known consequences that will be useful in the proof of our next result (Ross 2013, Asmussen 2003):
@@
    There exists a unique stationary distribution $\pi$ and the Markov chain is uniformly geometrically ergodic, i.e., there exists $K_{1}>0$ and $\rho\in[0,1)$ such that $\forall x\in\mathcal{S},$ we have
@@
    where $\bar{r}=E_{y\sim\pi}(r(y)),$ such that $||V(x)||\leq M_{V},\forall k\geq 0$ for some $M_{V}<\infty.$
@@
    Define $m_{k}=V(X_{k})-E(V(X_{k})|X_{k-1})$ for $k\geq 1.$ Clearly, $E(m_{k}|\mathcal{F}_{k-1})=0$ and $\{m_{k}\}_{k\geq 1}$ is a martingale difference sequence with respect to the filtration $\{\mathcal{F}_{k}\}_{k\geq 0}$ satisfying Assumption 1 with $\mathcal{F}_{k}=\sigma(X_{0},X_{1},X_{2},\ldots,X_{k})$. Clearly, from the previous bullet item, there exists constants $M_{2+\beta}$ and $M_{\beta}$ such that
  4. 4.
    While Theorem 1 holds for any $\Sigma_{\infty}\succ 0$, in this subsection we will find it convenient to define it to be
@@
    where $P_{ij}$ is the $(i,j)^{\rm th}$ element of the probability transition matrix. Due to the existence of a solution to Poisson’s equation, $\Sigma_{\infty}$ is well-defined.
@@
  1. 1.
    for any $\beta\in(0,1),$ $d_{\mathcal{W}}(U_{n},\Sigma_{\infty}^{1/2}Z)=\mathcal{O}\Bigl{(}\frac{1}{(1-\beta)n^{\beta/2}}\Bigr{)},$ and
  2. 2.
    $d_{\mathcal{W}}(U_{n},\Sigma_{\infty}^{1/2}Z)=\mathcal{O}\Bigl{(}\frac{\log n}{\sqrt{n}}\Bigr{)},$
@@
We start with the following standard use of Poisson’s equation and the definition of $m_{k}$ in Assumption 2 (see, for example, (Douc et al. 2018)) to write
@@
$U_{n}=W_{n}+\frac{V(X_{0})-V(X_{n})}{\sqrt{n}},\quad\mathrm{where}\quad W_{n}=\frac{1}{\sqrt{n}}\sum_{k=1}^{n}m_{k}.$ (7)
@@
The above expression goes to zero geometrically fast and thus,
@@
where $K_{3}$ is a problem-dependent constant.